# Demystifying Language Model Forgetting with Low-rank Example Associations

### v5

@@ Appendix E Towards Interpreting Fine-Grained Associations
We highlight some notable patterns in Table 6. (1) Some component $Z_{k}$ highlights forgetting patterns of upstream examples from certain domains. On OLMo-7B, the second component ($k=2$) highlights patterns where StackOverFlow examples are disproportionally less forgotten. (2) Some component $Z_{k}$ highlight forgetting when learning specific types of tasks. For example, the second component ($k=2$) on OLMo-1B highlights forgetting patterns after learning NLI tasks (mnli_matched, mnli_mismatched, snli). This also exemplifies how learning similar tasks cause a similar set of upstream examples to be more forgotten.
@@
Appendix F Synthetic Dataset Experiments
We mostly apply other training setups and hyperparameters in Aljundi et al. (2019a). Each task (rotation) includes 1,000 training examples. We train a MLP classifier to predict among 10 digits given an input image without providing its rotation (or the task identifier). We experiment MLP classification models with 1 layer (a linear model) to 5 layers. We collect the example associations $Z$, and visualize them in Figures 13 and 14. The upstream and the newly learned rotation tasks are ordered by their rotations in the $x$ or $y$ axis.
(a) Linear
(b) 2-Layer MLP
(c) 3-Layer MLP
(d) 4-Layer MLP
(e) 5-Layer MLP
Figure 13: Example associations between learned tasks and forgotten upstream examples on Rotated MNIST with overlapping ranges of rotations. We measure increase in the cross-entropy loss as forgetting.
(a) Linear
(b) 2-Layer MLP
(c) 3-Layer MLP
(d) 4-Layer MLP
(e) 5-Layer MLP
Figure 14: Example associations between learned tasks and forgotten upstream examples on Rotated MNIST with disjoint ranges of rotations.
@@ Appendix F Synthetic Dataset Experiments
MMLU abstract_algebra BBH boolean_expressions*
anatomy causal_judgement
astronomy date_understanding
business_ethics disambiguation_qa
clinical_knowledge dyck_languages*
college_biology* formal_fallacies*
college_chemistry geometric_shapes
college_computer_science hyperbaton*
college_mathematics logical_deduction_five_objects*
college_medicine* logical_deduction_seven_objects
college_physics logical_deduction_three_objects
computer_security movie_recommendation*
conceptual_physics* multistep_arithmetic_two
econometrics navigate
electrical_engineering object_counting*
elementary_mathematics penguins_in_a_table
formal_logic reasoning_about_colored_objects
global_facts* ruin_names
high_school_biology* salient_translation_error_detection
high_school_chemistry snarks
high_school_computer_science sports_understanding
high_school_european_history* temporal_sequences
high_school_geography tracking_shuffled_objects_five_objects
high_school_government_and_politics tracking_shuffled_objects_seven_objects
high_school_macroeconomics tracking_shuffled_objects_three_objects
high_school_mathematics web_of_lies
high_school_microeconomics word_sorting
high_school_physics* TruthfulQA Nutrition
high_school_psychology Stereotypes
high_school_statistics Confusion
high_school_us_history* Psychology
high_school_world_history Language
human_aging* Sociology
human_sexuality* Finance
international_law Indexical Error
jurisprudence Science
logical_fallacies* Misconceptions
machine_learning Economics
management* Education
marketing* Proverbs
medical_genetics Conspiracies
miscellaneous Religion
moral_disputes Statistics
moral_scenarios* Misquotations
nutrition Subjective
philosophy* Law
prehistory History
professional_accounting Fiction
professional_law Mandela Effect
professional_medicine* Politics
professional_psychology Misinformation
public_relations* Logical Falsehood
security_studies Distraction
sociology* Weather
us_foreign_policy* Myths and Fairytales
virology Superstitions
world_religions Advertising
Paranormal
Health

### v6

@@ Appendix G Towards Interpreting Fine-Grained Associations
We highlight some notable patterns in Table 8. (1) Some component $Z_{k}$ highlights forgetting patterns of upstream examples from certain domains. On OLMo-7B, the second component ($k=2$, also visualized in Figure 10(c)) highlights patterns where StackOverFlow examples are disproportionally more or less forgotten. (2) Some components $Z_{k}$ highlight forgetting when learning specific types of tasks. For example, the second component ($k=2$) on OLMo-1B highlights forgetting patterns after learning NLI tasks (mnli_matched, mnli_mismatched, snli). This also exemplifies how learning similar tasks causes similar sets of upstream examples to be more forgotten.
Appendix H Effect of Model Training Setups on Forgetting
In this section, we examine patterns of forgetting and its low-rank approximations across various training setups.
Table 9: $R^{2}$ of approximating forgetting association matrix $Z$ across different learning rates (LR) and batch sizes (BS). The default learning rate and batch sizes are 2e-6 and 8. We report results at $M=19$ learned tasks from Tulu and Dolly and $N=141,871$ upstream examples from Dolma.
Model Setting $r=2$ $r=3$ $r=5$
OLMo-7B Default 0.6981 0.7542 0.8232
LR=1e-6 0.7071 0.7697 0.8361
BS=32 0.8041 0.8453 0.9011
LR=5e-6 0.9331 0.9582 0.9718
LoRA, LR=1e-4 0.8963 0.9307 0.9562
OLMo-1B Default 0.8344 0.8836 0.9177
LR=1e-6 0.8471 0.8989 0.9214
BS=32 0.8724 0.9289 0.9607
Low-rank approximations hold at various learning rates, batch sizes, and LoRA fine-tuning. Table 9 summarizes $R^{2}$ of low-rank approximations. Reducing learning rate from the default 2e-6 to 1e-6 causes very minor change of $R^{2}$; Nevertheless, we notice increased $R^{2}$ under a larger learning rate like 5e-6, indicating even simpler associations between the learned tasks and forgotten upstream examples. Under this overly large learning rate, the model suffers substantially more forgetting on all examples. Besides, increasing the batch size (under the same number of parameter update steps) causes an increase in $R^{2}$, indicating simpler associations, possibly due to reduced variance of learned models under a larger batch size. LoRA fine-tuning Hu et al. (2022) results in simpler associations despite the model forgets less than full parameter fine-tuning.
Table 10: Sample Pearson correlation coefficient between upstream example forgetting under different replay strategies. The results are collected with OLMo-7B models fine-tuned on 8 Dolly tasks.
Setup A Setup B Sample Pearson $r$
No Replay No Replay + seed change 0.9403
No Replay Random Replay 0.7919
No Replay MF-Offline Replay 0.8121
Random Replay MF-Offline Replay 0.7709
Replay can change the patterns of forgetting compared to no replay, but the change is mild. We compute sample Pearson correlation coefficient between upstream example forgetting collected under different replay strategies, and more specifically (1) no replay (2) random replay and (3) MF-Offline replay on OLMo-7B and 8 Dolly tasks. We evaluate forgetting of upstream examples held-out from replay. We include correlations between upstream example forgetting under two different training random seeds as the baseline.
We summarize the results in Table 10. The correlations between no replay and replay are around 0.8, lower than the 0.94 baseline, but are still strongly positive. Replay example selection strategies (random or MF-offline) do not have a strong impact on the correlation. The results imply replay changes patterns of forgetting to a mild degree. Future works can study the limit where the patterns of forgetting experience notable changes.
Appendix I Combining Embedding and Matrix Completion based Prediction
Table 11: Combining matrix completion and embedding based prediction of forgetting. We report RMSE ($\downarrow$) or F1($\uparrow$) of predicting example forgetting over a held-out set of upstream examples after fine-tuning LMs on unseen new tasks.
In-Domain Out-of-Domain
Model OLMo-1B OLMo-7B MPT
OLMo-7B
Instruct
OLMo-1B OLMo-7B MPT
OLMo-7B
Instruct
Metrics RMSE RMSE RMSE RMSE F1 RMSE RMSE RMSE RMSE F1
Matrix Completion 2.79 7.14 10.41 13.74 58.16 2.82 5.76 7.03 38.90 43.57
Embedding 2.81 7.44 13.86 13.94 55.46 3.53 6.16 10.59 41.22 42.95
Embedding + MC 2.76 7.09 10.20 13.25 59.36 2.76 5.75 7.00 40.36 42.98
Although our results in Sec. 4 demonstrate subpar performance of the embedding-based approach, we show combining embedding and matrix completion approaches leads to improvement over the both. Specifically, we fit the residual error of the matrix completion model $Z-Z_{MC}$ with the embedding model, where $Z_{MC}$ is the prediction given by the matrix completion algorithms. We use the better of the additive and MF models reported in Table 3. Table 11 summarizes the results. We see that combining the two approaches improves performance over the two.
We consider the embedding based approaches better capture fine-grained knowledge conflicts that leads to forgetting of one example by learning the other (e.g., two documents stating contradictory facts) that is missed by matrix completion approaches that rely on statistics of forgetting.
Appendix J Synthetic Dataset Experiments
@@ Appendix I Combining Embedding and Matrix Completion based Prediction
We mostly apply other training setups and hyperparameters in Aljundi et al. (2019a). Each task (rotation) includes 1,000 training examples. We train a MLP classifier to predict among 10 digits given an input image without providing its rotation (or the task identifier). We experiment MLP classification models with 1 layer (a linear model) to 5 layers. We collect the example associations $Z$, and visualize them in Figure 12. The upstream and the newly learned rotation tasks are ordered by their rotations in the $x$ or $y$ axis.
(a) Linear, overlapping rotations
(b) Linear, disjoint rotations
(c) Linear, overlapping rotations
(d) Linear, disjoint rotations
(e) Linear, overlapping rotations
(f) Linear, disjoint rotations
(g) Linear, overlapping rotations
(h) Linear, disjoint rotations
(i) Linear, overlapping rotations
(j) Linear, disjoint rotations
Figure 12: Example associations between learned tasks and forgotten upstream examples on Rotated MNIST with overlapping or disjoint ranges of rotations. We measure increase in the cross-entropy loss as forgetting.
@@
MMLU abstract_algebra object_counting*
anatomy penguins_in_a_table
astronomy reasoning_about_colored_objects
business_ethics ruin_names
clinical_knowledge salient_translation_error_detection
college_biology* snarks
college_chemistry sports_understanding
college_computer_science temporal_sequences
college_mathematics tracking_shuffled_objects_five_objects
college_medicine* tracking_shuffled_objects_seven_objects
college_physics tracking_shuffled_objects_three_objects
computer_security web_of_lies
conceptual_physics* word_sorting
econometrics TruthfulQA Nutrition
electrical_engineering Stereotypes
elementary_mathematics Confusion
formal_logic Psychology
global_facts* Language
high_school_biology* Sociology
high_school_chemistry Finance
high_school_computer_science Indexical Error
high_school_european_history* Science
high_school_geography Misconceptions
high_school_government_and_politics Economics
high_school_macroeconomics Education
high_school_mathematics Proverbs
high_school_microeconomics Conspiracies
high_school_physics* Religion
high_school_psychology Statistics
high_school_statistics Misquotations
high_school_us_history* Subjective
high_school_world_history Law
human_aging* History
human_sexuality* Fiction
international_law Mandela Effect
jurisprudence Politics
logical_fallacies* Misinformation
machine_learning Logical Falsehood
management* Distraction
marketing* Weather
medical_genetics Myths and Fairytales
miscellaneous Superstitions
moral_disputes Advertising
moral_scenarios* Paranormal
nutrition Health
philosophy* Dolly brainstorming
prehistory rte closed_qa
professional_accounting snli information_extraction
professional_law wnli classification
professional_medicine* FLAN/Paraphrase glue_mrpc open_qa
professional_psychology glue_qqp* general_qa
public_relations* paws_wiki creative_writing
security_studies stsb summarization
sociology* OLMo2SFT-Mix coconot*
us_foreign_policy* evol_codealpaca_heval
virology flan_v2
world_religions no_robots
BBH boolean_expressions* numinamath_tir_math*
causal_judgement oasst1
date_understanding personahub_code
disambiguation_qa personahub_ifdata
dyck_languages* personahub_math*
formal_fallacies* aya
geometric_shapes open_math_2_gsm8k
hyperbaton* personahub_math_interm_algebra
logical_deduction_five_objects* sciriff
logical_deduction_seven_objects synthetic_finalresp_wildguardmixtrain
logical_deduction_three_objects table_gpt
movie_recommendation* wildchat
multistep_arithmetic_two wildjailbreak*
navigate personas-math-grade